OpenStudy (anonymous):

Suppose that y varies inversely with x. Write an equation for the inverse variation. y = 6 when x = 8 A.) Y = x/48 B.) Y = 48/x C.) X = Y/2 D.) Y = 2x

OpenStudy (anonymous):

@geerky42

OpenStudy (ali0094):

when u have options....you have to try them....you should put 9 instead of X and -12 instead of Y and then both side must be equal

OpenStudy (anonymous):

I think its C. I'm not sure

geerky42 (geerky42):

Can you double check your question? None of these options are correct.

OpenStudy (anonymous):

@geerky42 alright? I'll check

OpenStudy (dmndlife24):

(9, -12) 9 = x -12 = y -12 = 9/48 False -12 = 48/9 False 9 = -12/2 False -12 = 2(9) False None of them check so idk :/

OpenStudy (anonymous):

its X over 48 then 48 over X then Y over 2 then 2x

OpenStudy (anonymous):

@geerky42

OpenStudy (anonymous):

Oh i see the problem

OpenStudy (jdoe0001):

( 9 , 12 ) x y when x = 9, y = 12 test that in your choices, set x = 9 and see which one gives you y = 12 that one that does, is the solution

OpenStudy (anonymous):

the question was wrong

OpenStudy (anonymous):

@dan815